# Generative-Enhanced Heterogeneous Graph Contrastive Learning

### v2

@@ 4.4. Model Optimization
Node classification. First, the common practice for the node classification task (Park et al., 2020; Wang et al., 2021) is to train embeddings for all nodes using self-supervised loss. Then, we fine-tune the model on different datasets using a few labeled samples and a single-layer MLP. The node-level and edge-level strategies we propose typically lead to considerable loss differences in the initial phase. Therefore, it is necessary to balance their weights during the training phase. The loss generated by hierarchical contrastive learning is as follows:
(12) $\mathcal{L}_{hcl}=\lambda_{Intra}\cdot\mathcal{L}_{{Intra}}+\lambda_{Inter}\cdot\mathcal{L}_{{Inter}}$
where $\lambda_{{Intra}}$ and $\lambda_{{Inter}}$ are hyper-parameters to balance the influences of $\mathcal{L}_{{Intra}}$ and $\mathcal{L}_{{Inter}}$, respectively.
Link prediction. Next, for the link prediction task (Chen et al., 2023; Shi et al., 2018), we introduce the Bayesian Personalized Ranking Loss (BPR) (Rendle et al., 2012). This loss function treats two connected nodes as a positive pair and two unconnected nodes as a negative pair. Its objective is to maximize the scores between positive pairs and negative pairs. Formally, the BPR loss is defined as follows:
(13) $\mathcal{L}_{bpr}=-\frac{1}{|\mathcal{O}|}\sum_{(i,j,k)\in O}\log\sigma\left(d_{i}^{\top}d_{j}-d_{i}^{\top}d_{k}\right)$
where $\mathcal{O}=\left\{(i,j,k)\mid A_{i,j}=1,A_{i,k}=0\right\}$ is the training data. We combine $\mathcal{L}_{{hcl}}$ and $\mathcal{L}_{bpr}$ to get the link prediction loss.
(14) $\mathcal{L}=\lambda_{hcl}\cdot\mathcal{L}_{{hcl}}+\lambda_{bpr}\cdot\mathcal{L}_{{bpr}}$
@@
Time complexity analysis. The network schema view’s time complexity is $O\left(|V|\cdot D_{n}d^{2}\right)$, where $|V|$ denotes the edge number of a heterogeneous graph. $D_{n}$ is the average node degree and $d$ is the feature dimension. For the meta-path view, GC-HGNN has $M$ meta-path-based subgraphs. So, the time complexity of the meta-path view is $O\left(|\mathcal{V}|\cdot 2MLDd\right)$, where $L$ is the number of GNN layers. Then, the self-supervised loss contains $\lambda_{Intra}$ and $\lambda_{Inter}$. The time complexity to get positive samples and calculate $\lambda_{Intra}$ is $O\left(|\widetilde{\mathcal{V}}|\cdot d\right)$. For the $\lambda_{Inter}$, the complexity is $O\left(|\mathcal{V}|^{2}\cdot d\right)$. Because the edge set $\mathcal{E}$ is a matrix $\mathcal{E}\in\{0,1\}^{N\times N}$ composed of target nodes. In the link prediction task, the complexity of the additional $\mathcal{L}_{{bpr}}$ is $O\left(|\mathcal{O}|\cdot d\right)$. Thus, the overall time complexity of GC-HGNN is $O\left(|V|\cdot D_{n}d^{2}+|\mathcal{V}|\cdot 2MLD_{m}d+|\widetilde{\mathcal{V}}|\cdot d+|\mathcal{V}|^{2}\cdot d+|\mathcal{O}|\cdot d\right)$. This complexity is the time complexity for each epoch during training, and the number of training epochs will determine the overall training time.
5. EXPERIMENT
  • RQ1: How does GC-HGNN perform w.r.t. node classification task?
  • RQ2: How does GC-HGNN perform w.r.t. link prediction task?
  • RQ3: Are the key components in our GC-HGNN delivering the expected performance gains?
  • RQ4: How does GC-HGNN combine contrast and generation to achieve performance beyond one or the other?
  • RQ5: How do different settings influence the effectiveness of GC-HGNN?
Table 1. Statistics of datasets used in this paper.
@@ 5. EXPERIMENT
Table 2. Quantitative results (%) on node classification.
@@
5.1. Experimental Setings
5.1.1. Datasets
The performance of GC-HGNN has been validated on six real-world datasets. These datasets encompass four datasets for node classification: ACM, DBLP, Aminer, and Freebase, as well as four datasets for link prediction: Last.fm, Yelp, Douban Movie and Douban Book. A statistical overview of all the datasets is presented in Table 5.
@@ 5.1.1. Datasets
  • ACM (Han et al., 2022). The target nodes are papers (3 classes), averagely linked to 3.33 authors and one subject.
  • DBLP (Fu et al., 2020). The target nodes are authors (4 classes), averagely linked to 4.84 papers.
  • Freebase (Tian et al., 2023). The target nodes are movies (3 classes), averagely linked to 18.7 actors, 1.07 directors, and 1.83 writers.
  • Aminer (Wang et al., 2021). The target nodes are papers (4 classes), averagely linked to 2.74 authors and 8.96 references.
@@
  • Last.fm [^1] https://www.last.fm/ predicts user-artist interactions in a music recommendation setting, with node types including user, artist, and tag.
  • Yelp [^2] https://www.yelp.com/ predicts user-business interactions in a business recommendation setting, with node types including category, city, etc.
  • Douban Movie [^3] https://m.douban.com/ predicts user-movie interactions in a movie recommendation setting, with node types including actor, etc.
  • Douban Book (Shi et al., 2018) predicts user-book interactions in a book recommendation setting, with node types including author, etc.
5.1.2. Baselines
@@
  • DeepWalk (Perozzi et al., 2014) maps network structure to vector space via random walks, preserving neighbor similarity.
  • MP2vec (Dong et al., 2017) employs meta-path-guided random walks for heterogeneous network embeddings.
  • HERec (Shi et al., 2018) integrates meta-path-based random walks for node embeddings in HINs with fusion functions.
  • HAN (Wang et al., 2019) employs node-level and semantic-level attention to capture node and meta-path importance.
  • HGT (Hu et al., 2020a) introduces a novel architecture for handling heterogeneous structures in Web-scale graphs.
  • DMGI (Park et al., 2020) achieves self-supervised learning by meta-path-based networks and mutual infromation between networks.
  • HeCo (Wang et al., 2021) uses collaborative supervised contrast between network schema view and meta-path view.
  • MHGCN (Yu et al., 2022a) utilizes multi-layer convolution aggregation to capture meta-path interactions of different lengths.
  • HGCML (Wang et al., 2023) uses meta-paths to generate multi-views combined with comparative learning between views.
  • HGMAE (Tian et al., 2023) captures comprehensive graph information by reconstructing masked meta-path-based view and embedding.
@@
  • BPRMF (Rendle et al., 2012) learns implicit feature vectors between users and items to maximize difference in pairwise preferences.
  • LightGCN (He et al., 2020a) removes the feature transformation and nonlinear activations of GCN to achieve light recommendation.
  • SGL (Wu et al., 2021) generates contrast views by edge dropout to aid contrastive learning to enhance recommendation.
  • SMIN (Long et al., 2021) utilizes metagraph informax network to enhance user preference representation for social recommendation.
  • NCL (Lin et al., 2022) considers the relationship between neighbor nodes to enhance collaborative filtering.
  • HGCL (Chen et al., 2023) integrates heterogeneous relational semantics by leveraging contrastive self-supervised learning.
  • VGCL (Yang et al., 2023) introduces a variational graph reconstruction for graph augmentation to achieve light and robust recommendation.
5.2. Node Classification (RQ1)
The experiment results are shown in Table 5. Our experiments follow the settings of OpenHGNN (Han et al., 2022), a comprehensive and unified open-source platform. We categorize the baseline methods into three classes based on the input: unsupervised, supervised, and self-supervised. The symbols $\mathbf{X}$, $\mathbf{A}$, and $\mathbf{Y}$ in the table refer to the node features, graph structure, and predicted node labels, respectively.
Our GC-HGNN demonstrates state-of-the-art model (SOTA) performance on all datasets. Notably, self-supervised methods generally outperform shallow unsupervised methods. This is because unsupervised models only consider the graph structure, while self-supervised models attempt to mine hidden semantic relationships from the HIN. On the other hand, the performance of supervised models relies on more labels, which goes against the challenge of label scarcity and preciousness. HGMAE, as a generative reconstruction model, showcases the enormous potential of autoencoders in graph representation learning, and HeCo, as a contrastive model, also achieves remarkable results. These two models learn graph representations from different perspectives, providing ideas for designing subsequent graph neural network models. Inspired by this, GC-HGNN combines the generative masked autoencoder with a contrastive discriminator to enhance performance in contrastive learning by generating more hard negative samples.
5.3. Link Prediction (RQ2)
The recommendation has become a classic task for heterogeneous link prediction due to its wide application range. Considering the task adaptability, we apply LightGCN (He et al., 2020a) for the network schema. Here, we unify the embedding dimension to 64 and select the classic recall and NDCG (Zhang et al., 2023) as evaluation metrics. It is worth mentioning that our GC-HGNN compares with the latest general recommendation models.
The results are shown in Table 5.3. The first part presents the MF models, followed by general recommendation models, and the last part showcases specialized HGNN models. It can be observed that NCL and HGCL, as representatives of contrastive learning, demonstrate clear advantages over other models in their respective categories. The essence of contrastive learning lies in uniformity and alignment. However, modifying the graph structure to achieve uniformity is time-consuming and tricky, so it is better to consider this aspect from the perspective of the embedding space. Our GC-HGNN incorporates embeddings learned from the meta-path view into the network schema view and performs hierarchical contrastive learning. Such embeddings can better help the model learn a more uniform distribution because each meta-path brings different semantic information. Finally, across multiple datasets, our model achieved improvements of over 3% compared to the strongest baseline models, demonstrating the effectiveness of our model in aligning different meta-paths and perspectives.
Table 3. Quantitative results on link prediction.
@@ 5.3. Link Prediction (RQ2)
Figure 5. The comparison of GC-HGNN and its variants.
5.4. Ablation Study(RQ3)
In this subsection, we conducted ablation experiments to validate the effectiveness of key components in our GC-HGNN and made possible explanations regarding the experimental results. Here are three variants for our model.
  • $\boldsymbol{w}/\boldsymbol{o}{-}\mathbf{mae}$: This variant replaces the meta-path view’s masked autoencoder (MAE) with a normal GAT (Veličković et al., 2018) encoder.
@@ 5.4. Ablation Study(RQ3)
  • $\boldsymbol{w}/\boldsymbol{o}{-}\mathbf{hcl}$: The third cancels the hierarchical contrastive learning (HCL) module and directly uses the InfoNCE (2010infonce) loss for downstream tasks.
As shown in Figure 5, it can be observed that, in all cases, removing the HCL module results in worse performance compared to other variants. Additionally, variants where SAM and HCL are removed generally exhibit the worse performance in most cases. This further highlights two important points: 1) The graph structural information is overemphasized, and reconstructing the graph links is not necessary; 2) The existing heterogeneous contrastive learning strategies are not sufficient to serve as good contrastive discriminators to enhance the masked autoencoder effectively.
(a) MP2vec
(b) HeCo
(c) HGMAE
(d) GC-HGNN
Figure 6. Embedding visualization of DBLP dataset. Different colors indicate different node category labels.
(a) ACM
(b) DBLP
(c) Aminer
(d) Freebase
Figure 7. The impact wrt. hierarchical contrast coefficient.
@@
Figure 8. The impact wrt. dimension and mask ratio. The axes on the left belong to the ACM and DBLP datasets and the axes on the right belong to other datasets.
5.5. Interpretability and Visualization (RQ4)
In this study, we focus on discussing the various modules of GC-HGNN, attempting to provide a rational concept or explanation. We investigated the advantages of the generative-contrastive model and then further illustrated the motivation through the representation visualization.
Interpretability. Generative-contrastive representation learning (Wu et al., 2023; Yang et al., 2023; Goodfellow et al., 2020) utilizes discriminative loss functions as objectives, aiming to address some inherent shortcomings of generative reconstruction, such as sensitive and conservative distribution and low-level abstraction objectives. In contrastive learning (Chen et al., 2020; Wang et al., 2021; He et al., 2020b; Wu et al., 2021), we can alleviate these issues by setting up appropriate contrastive discriminators to learn ‘distinguishable’ information for different samples. However, the encoder and decoder brought by generative learning (Kenton and Toutanova, 2019; Hou et al., 2022; Tian et al., 2023; Chowdhery et al., 2023) endow it with powerful embedding expression capabilities. The decoder requires the representation to be reconstructable, meaning that it contains all the information of the input. Furthermore, generative-contrastive self-supervised methods abandon the point-to-point reconstruction objective and instead pursue the alignment of distributions in the data. In summary, adversarial methods (Liu et al., 2021) absorb the advantages of both generative and contrastive methods, but also face the challenge of how to build a framework in the field of heterogeneous graphs. We propose a masked autoencoder to enhance the expressive power of meta-path perspectives, and then use hierarchical contrastive learning and designed sampling strategies to serve as enhanced contrastive discriminators. Experiments demonstrate that this is an effective and novel exploration.

### v3

@@ III-D Model Optimization
Node classification. First, the common practice for the node classification task [45, 7] is to train embeddings for all nodes using self-supervised loss. Then, we fine-tune the model on different datasets using a few labeled samples and a single-layer MLP. The node-level and edge-level strategies we propose typically lead to considerable loss differences. Therefore, it is necessary to balance their weights during the training phase. The loss generated by hierarchical contrastive learning is as follows:
$\mathcal{L}_{hcl}=\lambda_{Intra}\cdot\mathcal{L}_{{Intra}}+\lambda_{Inter}\cdot\mathcal{L}_{{Inter}}+\lambda_{Gen}\cdot\mathcal{L}_{{Gen}}$ (13)
where $\lambda_{{Intra}}$, $\lambda_{{Inter}}$ and $\lambda_{Gen}$ are hyper-parameters to balance the influences of losses.
Link prediction. Next, for the link prediction task [5, 4], we introduce the Bayesian Personalized Ranking Loss (BPR) [46]. This loss function treats two connected nodes as a positive pair and two unconnected nodes as a negative pair. Its objective is to maximize the scores between positive pairs and negative pairs. Formally, the BPR loss is defined as follows:
$\mathcal{L}_{bpr}=-\frac{1}{|\mathcal{O}|}\sum_{(i,j,k)\in O}\log\sigma\left(\mathbf{d}_{i}^{\top}\mathbf{d}_{j}-\mathbf{d}_{i}^{\top}\mathbf{d}_{k}\right)$ (14)
where $\mathcal{O}=\left\{(i,j,k)\mid A_{i,j}=1,A_{i,k}=0\right\}$ is the training data. We combine $\mathcal{L}_{{hcl}}$ and $\mathcal{L}_{bpr}$ to get the loss:
$\mathcal{L}=\lambda_{hcl}\cdot\mathcal{L}_{{hcl}}+\lambda_{bpr}\cdot\mathcal{L}_{{bpr}}$ (15)
@@
III-E Time Complexity Analysis
The network schema view’s time complexity is $O\left(|V|\cdot D_{n}d^{2}\right)$, where $|V|$ denotes the edge number of a heterogeneous graph. $D_{n}$ is the average node degree and $d$ is the feature dimension. For the meta-path view, GC-HGNN has $M$ meta-path-based subgraphs. So, the time complexity of the meta-path view is $O\left(|\mathcal{V}|\cdot 2MLDd\right)$, where $L$ is the number of GNN layers. Then, the self-supervised loss contains $\lambda_{Intra}$ and $\lambda_{Inter}$. The time complexity to get positive samples and calculate $\lambda_{Intra}$ is $O\left(|\widetilde{\mathcal{V}}|\cdot d\right)$. For the $\lambda_{Inter}$, the complexity is $O\left(|\mathcal{V}|^{2}\cdot d\right)$. Because the edge set $\mathcal{E}$ is a matrix $\mathcal{E}\in\{0,1\}^{N\times N}$ composed of target nodes. In the link prediction task, the complexity of the additional $\mathcal{L}_{{bpr}}$ is $O\left(|\mathcal{O}|\cdot d\right)$. Thus, the overall time complexity of GC-HGNN is $O\left(\left(|V|\cdot D_{n}d+|\mathcal{V}|\cdot 2MLD_{m}+|\widetilde{\mathcal{V}}|+|\mathcal{V}|^{2}+|\mathcal{O}|\right)\cdot d\right)$. This complexity is the time complexity for each epoch during training, and the number of training epochs will determine the overall training time. We compared several strong baselines in terms of training time. Our model performs faster per epoch due to optimized modules, demonstrating lower complexity and greater scalability compared to baseline methods.
TABLE II: Time and resource comparison on datasets.
{NiceTabular}l c c c c Dataset ACM DBLP Aminer Freebase
Time (eopch/s)
HeCo [7] 0.204 0.198 0.361 0.272
HGMAE [23] 0.212 0.334 0.203 0.114
GC-HGNN 0.033 0.058 0.043 0.037
Resource (memory/G)
HeCo [7] 2.21 3.25 8.75 6.36
HGMAE [23] 5.49 6.82 14.48 8.93
GC-HGNN 2.54 4.10 8.33 6.65
TABLE III: Statistics of datasets used in this paper.
@@ III-E Time Complexity Analysis
IV EXPERIMENT
  • RQ1: How does GC-HGNN perform w.r.t. tasks of node classification and link prediction?
  • RQ2: Are the key components in our GC-HGNN delivering the expected performance gains?
  • RQ3: How does GC-HGNN combine contrast and generation to achieve performance beyond one or the other?
  • RQ4: How do different settings influence the effectiveness of GC-HGNN?
TABLE IV: Quantitative results (%) on node classification. In the table, the symbols $\mathbf{X}$, $\mathbf{A}$, and $\mathbf{Y}$ denote the node features, graph structure, and predicted node labels, respectively.
@@ IV EXPERIMENT
IV-A Experimental Setings
We followed the experimental benchmarks of HeCo and HGMAE, where the hidden layer dimension is fixed at 128, and the number of fine-tuning samples per class is set to 20, 40, and 60. The original HGMAE paper used a hidden layer dimension of 512, for which we readjusted the optimal parameters. Due to space limitations, we reported the results for 20 samples, as labels are usually valuable, and fewer samples often require stronger model generalization ability. All experimental results were repeated 10 times on the 4090, and we release the reproducible framework for the model as open source at the code [^1] https://github.com/wangyu0627/HGNN-baselines.
IV-B Datasets
The performance of GC-HGNN has been validated on six real-world datasets. These datasets encompass four datasets for node classification: ACM, DBLP, Aminer, and Freebase, as well as four datasets for link prediction: Last.fm, Yelp, Douban Movie and Douban Book. A statistical overview of all the datasets is presented in Table III-E.
@@ IV-B Datasets
  • ACM [47]. The target nodes are papers (3 classes), averagely linked to 3.33 authors and one subject.
  • DBLP [33]. The target nodes are authors (4 classes), averagely linked to 4.84 papers.
  • Freebase [23]. The target nodes are movies (3 classes), which are typically linked to an average of 18.7 actors, 1.07 directors, and 1.83 writers.
  • Aminer [7]. The target nodes are papers (4 classes), averagely linked to 2.74 authors and 8.96 references.
@@
  • Last.fm [^2] https://www.last.fm/ predicts user-artist interactions in a music recommendation setting, with other node type including tag.
  • Yelp [^3] https://www.yelp.com/ predicts user-business interactions in a business recommendation setting, with node types including category, city and category,etc.
  • Douban Movie [^4] https://m.douban.com/ predicts user-movie interactions in a movie recommendation setting, with node types including actor, director and type, etc.
  • Douban Book [4] predicts user-book interactions in a book recommendation setting, with node types including author, publisher and group, etc.
IV-C Baselines
@@
  • DeepWalk [48] maps network structure to vector space via random walks, preserving neighbor similarity.
  • MP2vec [3] employs meta-path-guided random walks for heterogeneous network embeddings.
  • HERec [4] integrates meta-path-based random walks for node embeddings in HINs with fusion functions.
  • HAN [2] employs node-level and semantic-level attention to capture node and meta-path importance.
  • HGT [36] introduces a novel architecture for handling heterogeneous structures in Web-scale graphs.
  • DMGI [45] achieves self-supervised learning by meta-path-based networks and mutual infromation between networks.
  • HeCo [7] uses collaborative supervised contrast between network schema view and meta-path view.
  • MHGCN [49] utilizes multi-layer convolution aggregation to capture meta-path interactions of different lengths.
  • HGCML [17] uses meta-paths to generate multi-views combined with comparative learning between views.
  • HGMAE [23] captures graph information by reconstructing masked meta-path-based view and embedding.
@@
  • BPRMF [46] learns implicit feature vectors between users and items to maximize difference in pairwise preferences.
  • LightGCN [30] removes the feature transformation and nonlinear activations of GCN to achieve recommendation.
  • SGL [15] generates contrast views by edge dropout to aid contrastive learning to enhance recommendation.
  • SMIN [6] utilizes metagraph informax network to enhance user preference representation for social recommendation.
  • NCL [20] considers the relationship between neighbor nodes to enhance collaborative filtering.
  • HGCL [5] integrates heterogeneous relational semantics by leveraging contrastive self-supervised learning.
  • VGCL [16] introduces a variational graph reconstruction for graph augmentation to achieve robust recommendation.
TABLE V: Quantitative results on link prediction.
@@
IV-D Overall Performance (RQ1)
Node Classification. The experiment results are shown in Table IV. Notably, self-supervised methods generally outperform unsupervised ones based on shallow networks. This performance disparity arises because unsupervised models only consider the graph structure, whereas self-supervised models endeavor to mine hidden semantic relationships from the HIN. Conversely, the performance of supervised models is dependent on more labels, which goes against the challenge of label scarcity and preciousness. HGMAE, a generative reconstruction model, demonstrates the significant potential of autoencoders in graph representation learning, while HeCo, a contrastive model, also yields remarkable results. These two models learn graph representations from different perspectives, providing insights for designing future GNN models. Inspired by these insights, GC-HGNN integrates the generative masked autoencoder with a contrastive discriminator to improve performance by generating more hard negative samples.
Link Prediction. Recommendation is a classic task for heterogeneous link prediction. Considering the task adaptability, we use LightGCN [30] to replace the network schema. Here, we standardize the embedding dimension to 64 and select the recall and NDCG [50] as evaluation metrics. The results are shown in Table IV-C. It can be observed that NCL and HGCL, as representatives of contrastive learning, demonstrate distinct advantages over other models within their respective categories. The essence of contrastive learning lies in achieving uniformity and alignment. However, modifying the graph structure to achieve uniformity is time-consuming and tricky, so it is better to consider this aspect from the perspective of the embedding space. Our GC-HGNN incorporates embeddings learned from the meta-path view into the network schema view and performs hierarchical contrastive learning. Finally, across four datasets, our model achieved improvements of over 3% compared to the strongest baseline models.
Figure 5: The comparison of GC-HGNN and its variants.
(a) MP2vec
(b) HeCo
(c) HGMAE
(d) GC-HGNN
Figure 6: Embedding visualization of DBLP dataset. Different colors indicate different node category labels.
@@ IV-D Overall Performance (RQ1)
Figure 7: The impact wrt. hierarchical contrast coefficient. The axes on the left belong to the ACM and DBLP datasets, and the axes on the right belong to other datasets.
IV-E Ablation Study(RQ2)
In this section, we conducted ablation experiments to validate the effectiveness of key components in our GC-HGNN and provided possible explanations for the experimental results. We present three variants of our model.
  • $\boldsymbol{w}/\boldsymbol{o}{-}\mathbf{MAE}$: This variant replaces the meta-path view’s masked autoencoder (MAE) with a normal GAT [29] encoder. Because usually GAT has stronger performance than GCN [30] on node classification tasks.
@@ IV-E Ablation Study(RQ2)
  • $\boldsymbol{w}/\boldsymbol{o}{-}\mathbf{INTRA}$: This variant replaces the intra-contrast corresponds to Eq. 10.
  • $\boldsymbol{w}/\boldsymbol{o}{-}\mathbf{INTER}$: This variant cancels the inter-contrast integrated from Eqs. 11 and 12.
  • $\boldsymbol{w}/\boldsymbol{o}{-}\mathbf{HCL}$: The last cancels the hierarchical contrastive learning (HCL) module and directly uses the InfoNCE [39] loss for downstream tasks.
As shown in Figure 5, in all cases, removing the HCL module results in poorest performance compared to other variants. Additionally, variants where SAM or HCL is removed generally exhibit the worse performance in most cases. This suggests that inter-contrast through aligned local and global heterogeneous information significantly enhances semantics, while intra-contrast regulates semantic uniformity. We further highlights two important points: 1) The graph structural information is overemphasized, and reconstructing the graph links is not necessary; 2) The existing heterogeneous contrastive learning strategies are insufficient to serve as effective contrastive discriminators to enhance the masked autoencoder. It is noteworthy that ‘INTER’ serves as the principal loss function. This suggests that aligning high-order and first-order semantic representations of heterogeneous information can significantly improve the prediction accuracy, while contrast between meta-paths regulates semantic consistency [41].
IV-F Interpretability and Visualization (RQ3)
In this study, we examine the various modules of GC-HGNN, attempting to provide a rational explanation. We investigated the advantages of the generative-contrastive model and then further illustrated the motivation through the representation visualization.
IV-F1 Interpretability
Generative-contrastive (GC) representation learning [19, 16, 18] employs discriminative loss functions as its objectives, aiming to address some inherent shortcomings of generative reconstruction, such as sensitive and conservative distribution issues and low-level abstraction objectives. In contrastive learning [39, 7, 9, 15], we can alleviate these issues by setting up appropriate contrastive discriminators to learn distinguishable information for different samples. However, the encoder and decoder in generative learning [12, 28, 23, 11] provide powerful embedding expression capabilities. The decoder requires the representation to be reconstructable, meaning that it contains all the information of the input. Furthermore, GC-based self-supervised methods eschew the point-to-point reconstruction objective and aim to align distributions in the data. In summary, These methods [14] integrate the advantages of both generative and contrastive methods, yet confront the challenge of constructing a framework within the field of heterogeneous graphs. We propose a masked autoencoder to enhance the expressive power of meta-path perspectives, followed by the use of hierarchical contrastive learning and sampling strategies to serve as enhanced contrastive discriminators. Our experiments demonstrate that this is an effective and novel exploration.
@@ IV-F1 Interpretability
Figure 8: The impact wrt. dimension and mask ratio. The axes on the left belong to the ACM and DBLP datasets, and the axes on the right belong to other datasets.
TABLE VI: Performance comparison with coefficient $\mathcal{L}_{Gen}$.
{NiceTabular}ccccccccc Dataset ACM DBLP Aminer Freebase
Lambda Ma-F1 Mi-F1 Ma-F1 Mi-F1 Ma-F1 Mi-F1 Ma-F1 Mi-F1
0.1 90.54$\pm$0.28 90.36$\pm$0.31 92.54$\pm$0.36 93.01$\pm$0.30 72.44$\pm$0.30 80.28$\pm$0.34 61.47$\pm$0.37 64.46$\pm$0.43
0.2 89.50$\pm$0.32 89.48$\pm$0.29 91.37$\pm$0.41 92.24$\pm$0.34 73.25$\pm$0.20 81.45$\pm$0.32 62.40$\pm$0.42 65.85$\pm$0.35
0.3 88.47$\pm$0.33 87.72$\pm$0.29 90.23$\pm$0.40 90.92$\pm$0.34 71.34$\pm$0.18 79.16$\pm$0.26 60.62$\pm$0.47 63.74$\pm$0.41
0.4 86.66$\pm$0.29 86.20$\pm$0.32 87.33$\pm$0.37 87.29$\pm$0.31 66.21$\pm$0.21 73.35$\pm$0.31 57.47$\pm$0.43 60.64$\pm$0.36
0.5 80.51$\pm$0.32 80.25$\pm$0.38 82.43$\pm$0.29 82.98$\pm$0.33 63.41$\pm$0.25 71.18$\pm$0.27 55.38$\pm$0.30 59.14$\pm$0.47
TABLE VII: Performance comparison of different encoder and decoder variants.
{NiceTabular}c—c—cccccccc Encoder Decoder ACM DBLP Aminer Freebase
GAT GAT $\underline{90.29\pm 0.22}$ $\underline{90.16\pm 0.30}$ $\underline{92.14\pm 0.32}$ $\underline{92.49\pm 0.41}$ $\mathbf{73.44}\pm\mathbf{0.20}$ $\mathbf{81.28}\pm\mathbf{0.32}$ $\mathbf{62.47}\pm\mathbf{0.42}$ $\mathbf{65.85}\pm\mathbf{0.35}$
GCN GCN $\mathbf{90.54}\pm\mathbf{0.28}$ $\mathbf{90.36}\pm\mathbf{0.31}$ $91.16\pm 0.24$ $91.52\pm 0.27$ $71.90\pm 0.32$ $79.48\pm 0.43$ $61.85\pm 0.20$ $64.88\pm 0.24$
GAT GCN $89.24\pm 0.27$ $89.46\pm 0.32$ $90.74\pm 0.18$ $91.15\pm 0.23$ $\underline{72.69\pm 0.25}$ $\underline{80.44\pm 0.34}$ $61.23\pm 0.33$ $63.78\pm 0.26$
GCN GAT $89.68\pm 0.18$ $89.61\pm 0.36$ $\mathbf{92.54}\pm\mathbf{0.36}$ $\mathbf{93.01}\pm\mathbf{0.30}$ $72.31\pm 0.44$ $79.98\pm 0.21$ $\underline{62.02\pm 0.29}$ $\underline{65.24\pm 0.35}$
IV-F2 Visualization